OpenStudy (anonymous):

PLEASE HELP!! WILL FAN AND GIVE MEDAL!! Solve the following inequality:

OpenStudy (anonymous):

\[3(2a-8)\le14a +6\]

Mehek (mehek14):

first distribute \(3(2a-8)\\ 3*2a = 6a\\3*-8=-24\\6a-24≤14a+6\) next combine like terms \(-30≤8a\)

OpenStudy (anonymous):

ok do i divide on both sides after that?

Mehek (mehek14):

what are your options?

OpenStudy (anonymous):

i read what u did but i think something is wrong

Mehek (mehek14):

no do you have any options?

Mehek (mehek14):

you just divide -30/8 = -3.75

OpenStudy (anonymous):

you said that -24 +6=-30 i dont think thats right

OpenStudy (anonymous):

wouldnt it be -17

Mehek (mehek14):

i subtracted -24 - 6 when you move something on the other side, you do the opposite opposite of + is -

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

ok nvm thanks anyway!

Mehek (mehek14):

yw :)